OpenStudy (anonymous):

A DVD rotates through an angle of 20π radians in a second. At this speed how many revolutions does the DVD make in 30 seconds?

jimthompson5910 (jim_thompson5910):

(20pi radians/1 second)*(30 seconds) = (20pi*30) radians (20pi radians/1 second)*(30 seconds) = 600pi radians So in 30 seconds, it rotates 600pi radians 2pi radians = 1 rev which allows us to say (600pi radians)*(1 rev/2pi radians) = (600pi/2pi) rev (600pi radians)*(1 rev/2pi radians) = 300 rev So in 30 seconds, it completes 300 full revolutions